Mr. Takahashi has n = 16 and he is trying to compute a 90% confidence interval. However, he does not have access to the population’s standard deviation and instead he is using a sample standard deviation. Mr. Takahashi should use a _-stat and it should be ___.

A. Z; 1.753

B. t; 1.753

C. t; 1.746

D. Z; 1.746
Answer #1

solution

Degrees of freedom = df = n - 1 = 16 - 1 = 15

At 90% confidence level the t is ,

= 1 - 90% = 1 - 0.90 = 0.1

/ 2 = 0.1 / 2 = 0.05

t /2,df = t0.05,15 = 1.753    ( using student t table)

Mr. Takahashi should use a _t -stat and it should be . t; 1.753___.